solve each system of equations by using cramers rule:
{3x - 4y + 2z = 1
{x - y + 2z = -2
{2x + 2y + 2z = -3

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

solve each system of equations by using cramers rule:
{3x - 4y + 2z = 1
{x - y + 2z = -2
{2x + 2y + 2z = -3
-------------
Step 1, find the determinant.
---

Answer by ikleyn(52781)

.
.
On Cramer's rule for solving systems of 3 equations in 3 unknowns see the lessons
    - Determinant of a 3x3 matrix
    - Co-factoring the determinant of a 3x3 matrix
    - HOW TO solve system of linear equations in three unknowns using determinant (Cramer's rule)
    - Solving systems of linear equations in three unknowns using determinant (Cramer's rule)
    - Solving word problems by reducing to systems of linear equations in three unknowns
in this site.

Also, you have this free of charge online textbook in ALGEBRA-II in this site
    - ALGEBRA-II - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic
     "3x3-Matrices, determinants, Cramer's rule for systems in three unknowns"